# RATSF: Empowering Customer Service Volume Management through Retrieval-Augmented Time-Series Forecasting

### v1

@@ 3.1 RACA
3.2 TSKB
In the development of TSKB, we employ an innovative strategy for constructing the indexing sequence (K) and content sequence (V). Unlike traditional methods that solely store and retrieve the entire original sequence, TSKB preserves the original sequence as the core content while selects certain segments from it to construct a discriminative indexing sequence. This approach can be likened to processing a written piece where conventional methods involve direct full-text searches for desired information, which may be inefficient and less precise; whereas with TSKB, it’s akin to extracting key headlines from the body of the text to serve as indices that facilitate rapid access to core information. This dual-sequence structure allows the system, when handling large-scale data, to efficiently locate and access targeted portions of the original sequence through customized index sequences, thereby significantly enhancing overall performance.
In this article, to illustrate our system design, we employ a fixed slicing logic to select segments from V for computing the retrieval representation. In section 3.3, we will present in detail how the representation of V is learned.
Figure 3: Within the cold start stage, we use raw sequence of retrieval input to find Top-K relevant sequences using DTW. After one epoch, we switch to Euclidean Distance for retrieval embedding to gather Top-K matches.
Figure 4: TSKB utilizes a rolling window of length $L_{v}$ to collect V, with an indexing segment of length $L_{r}$ taken from its leading part as K , and advances the window in steps of size S.
3.3 Learning Retrieval Embedding
Judging the quality of a retrieval representation mainly depends on how well it captures similarities at key information points relevant to forecasting tasks. Time series data is complex due to numerous information points and the challenge of presetting comparative weights. Therefore, we use this principle: if a representation closely mirrors the important details needed for prediction, its overall performance will be better. Later, we’ll explain the exact steps to put this principle into practice. We acquire the retrieval embedding for TSKB through two procedures.
3.3.1 Sequential Slicing
To obtain content sequence V, we utilize a rolling window approach where the step size is $S$ and the window length is $L_{v}$, to sample all historical sequences and incorporate them into TSKB. Since prediction models only use present data to infer the future, we align indexing sequence K with this constraint. We extract the initial segment of each V, having a length of $L_{r}$, to serve as its K. The selection of $L_{v}$, $L_{r}$ and $S$ is introduced in section 4.
3.3.2 Embedding Learning
To guarantee the embedding in capturing the essential information for forecasting, we utilize the encoder of the RATSF’s forecasting model. The model inherently excels at selecting precise information during training to improve forecast accuracy. In practice, every indexing sequence K from the knowledge base undergoes the RATSF forecasting model’s encoder layer to generate a retrieve embedding.
Meanwhile, in the early stages of training the forecasting model, its encoder is not yet capable of producing representations that accurately match target sequences. The effectiveness and learning pace of the RATSF’s forecasting model are significantly correlated with whether or not it retrieves historical sequences that are beneficial to the forecasting task.
To avoid time and data-consuming iterations stemming from a random initialization state, we introduce DTW as an auxiliary tool during the initial phase of model training, as shown in left part of Figure 3. Initially, DTW is utilized for similarity-based sequence retrieval, which is then incorporated into the iterative training process of the RATSF forecasting model. After finish one epoch of training, we transition to using embedding generated by the forecasting model itself for similar sequence retrieval, continuing the training until the model converges.
Figure 5: The lower plot displays the top 3 relevant retrieved sequences for one sample that are end-to-end connected. Each retrieved sequence has a length of 21 ($L_{f}$), comprising 14 points preceding the forecasting date and 7 points to be forecasted, highlighted with an orange background. The color bars represent aggregated cross-attention values at the first forecasting point, in which lighter color indicate higher attention value. In the upper part of the figure, predicted (yellow) and ground truth (orange) values align at corresponding positions with retrieved sequence. For this sample, the model notably focuses on segments indicating ascent and predicts a subsequent increase in value.
@@ 3.3.2 Embedding Learning
Table 1: Performance comparison of four Transformer models with and without the application of RATSF on three experimental datasets. All results are evaluated using MSE (Mean Squared Error) and MAE (Mean Absolute Error), and the best result is bolded.
@@
4 Experiments
In order to evaluate the generalizability and efficacy of RATSF in industry applications, we initially carry out a range of experiments to contrast the performance changes among typical time-series models when augmented with RATSF across various tasks, focusing particularly on the Fliggy Hotel Service Volume Dataset (FHSV). Following this, we employ a sequence of ablation studies to separately demonstrate the design choice of each improvement – RACA, TSKB, and Retrieval Embedding.
4.1 Experiment Setup
Datasets. Given that our primary focus lies in the forecasting task within the context of customer service volume management, we initially employed the Fliggy Hotel Service Volume Dataset to substantiate the efficacy of RATSF, the detail description of the dataset is in Appendix A. Additionally, to demonstrate the performance of our approach in other contexts, we employed two classic datasets: ETT and Exchange.
Models.We contrasted the vanilla Transformer against three of its prominent time-series domain variants: Fedformer, Autoformer, and NS-Transformer. Specifically, in keeping with the intrinsic sequential decomposition characteristics of Fedformer and Autoformer, historical segments were initially broken down using each model’s native calculations, for example the seasonal init and trend-cyclical init in Autoformer. Subsequently, these decomposed segments were embedded and fed into the RACA module for further processing.
Forecasting Setting.Consistent with the original intent of validating the effectiveness of RATSF in the context of customer service volume management, we adopt the business setting from Fliggy. In the Fliggy’s practical operations, we are required to finalize our forecasting for the upcoming week’s day-by-day staffing needs one week in advance, which is why the actual prediction window period is set at 7 days, represented as 7 tokens. In the travel industry, which Fliggy App’s serving, a quarter generally defines a relatively complete business trend. Therefore, we choose an encoder window length of 98 days—exceeding 90 days and divisible by 14—which consequently results in $\mathbf{X_{o}}$ is 98 units long. And we concate $\left\{x_{t-14},...,x_{t}\right\}$ with a placeholder of length 7 to form $\mathbf{X_{f}}$. To match the length of $\mathbf{X_{f}}$, we set $L_{v}$=21 and $L_{r}$=14. Regarding normalization, we employ unified $\mu$ and $\sigma$ values for whitening operations during the preprocessing stage and inverse normalization is carried out after forecasting. We have applied the same data processing scheme to both the ETT and Exchange datasets.
4.2 Main Restult
Table 1 presents the performance comparison of four Transformer models with and without the application of RATSF on three experimental datasets. In the scenario of forecasting with the Fliggy Hotel Customer Service Volume Dataset, the Transformer model’s MAE loss dropped by 18% after incorporating RATSF, potentially translating to a reduction of approximately 200 redundant personnel cost in practical management operations. Notably, even with the NSTransformer [17], Autoformer, and FedFormer models, which have undergone temporal optimization in their attention operators, adopted the RATSF design, they respectively recorded decreases in MAE loss by 14%, 5%, and 4%.
@@ 4.2 Main Restult
Additionally, across the ETT and Exchange datasets, the adoption of RATSF design in the Transformer and its time-series optimized variants led to noticeable reductions in MSE metrics. Moreover, among these, the NSTransformer consistently outperforms the other models, likely because its mean-adapted feature proves effective in a broad range of time-series prediction domains.
4.3 Ablation Study
Section 4.3.1 comparatively analyzes the differences in various designs for integrating historical pieces into forecasting. Section 4.3.2 examines the impact of integrating varying numbers of historical pieces on forecast effectiveness. Section 4.3.3 demonstrates the advantages of utilizing model encoder for recalling historical pieces.
Table 2: Ablation of the specifically designed structures within RACA on FHSV scenario.
@@ 4.3 Ablation Study
4.3.1 RACA’s design choice
@@
A deeper examination disclosed that the Cross Attention mechanism in the RACA design efficiently guides the forecast sequence to concentrate on past trends analogous to the current scenario. Figure 5 illustrates the attention map within the RACA for the data points under forecast, where highlighted color blocks indicate that the data point being forecast places significant attention on historical segments that exhibiting similar trend patterns, and these historical segments’ trends closely align with the actual waveform following the data point in the ground truth.
4.3.2 Optimal Retrieval Count
Table 3: Results of recalling different numbers of pieces in our RATSF model.
dataset methods metric
mse 4344450.512
FHSV baseline mae 1349.146
mse 3678230.020
recalled Top-1 pieces mae 1160.101
mse 3700934.754
recalled Top-2 pieces mae 1155.952
mse 3421546.027
recalled Top-3 pieces mae 1101.172
mse 3460814.538
recalled Top-4 pieces mae 1108.895
mse 3750426.751
recalled Top-5 pieces mae 1231.564
Having established that integrating effective historical pieces positively impacts forecasting, we then inquire about the optimal number of sequences to retrieve. To demonstrate the effect of this choice, we gradually increase the number of integrated retrieve sequences from 0 to 5. As shown in Table 3, during the process of expanding from integrating no historical piece to integrating up to three historical pieces, the MAE error consistently decreases; however, when the quantity of integrated historical pieces is further increased, the forecasting error starts to escalate.
Upon analyzing samples, the insight is obvious: not all of the top five retrieved historical piece for most of samples closely resemble the Ground Truth. The reason being, focusing solely on the TopK ranking without adequately considering similarity thresholds can easily lead to a scenario where, as K increases, the quality of the retrieved pieces becomes less assured and more prone to introduce noise and irrelevant information, thus potentially causing confusion instead.
4.3.3 Retrieval Embedding
We claimed that DTW is not optimal as the ultimate retrieving method. In this part of ablation we show that with an experiment on FHSV. We compare DTW, a trained MLP and current design choice, the forecasting model’s encoder in terms of accuracy.
Table 4 clearly demonstrates that the forecasting model’s encoder output provide embedding outperforms both the other two schemes significantly. The advantages of encoder output over MLP are evident, particularly because the attention mechanism within the transformer’s encoder is better suited to discern and refine the interdependence’s among data points in sequences, thereby abstracting more accurate patterns.
Table 4: Comparison of several variants of retrieval representation schemes.
dataset methods metric
@@ 4.3.3 Retrieval Embedding
FHSV baseline + encoder mae 1101.172
5 Conclusion
In this paper, we propose the RATSF and explain its effectiveness in enhancing time-series forecasting performance. Specifically, our approach improves from three dimensions: first, we integrate the RACA mechanism into the transformer model architecture, which leverages historical data segments similar to current contexts to bolster forecasting prowess. Second, we construct a KV schema for a time-series knowledge base that significantly enhances the flexibility and precision of the retrieval process. And third, we utilize the Encoder component within the forecasting model to embed K sequences, thereby refining the accuracy of retrieving relevant historical information. The method introduced in this study has been actively operational within Fliggy’s service volume management system for several months. Of particular importance is that due to the strong flexibility and scalability inherent in this design, the method can play a significant role across a wide array of practical application scenarios.
Appendix A Fliggy Hotel Service Volume Dataset
The Fliggy Hotel Service Volume Dataset (FHSV) is derived from service records within the Fliggy APP’s customer service center, capturing the volume of daily customer service operations related to inquiries about hotel reservations, cancellations, and changes. Each instance of a service request received either through the APP interface or by phone is counted as a single service event. Daily aggregates of such events are recorded, forming individual data points. This dataset comprises 1740 data points collected between January 1, 2019, and October 7, 2023.
To facilitate experimentation, this time series dataset has been sequentially partitioned in a chronological order into three subsets: the train set (from January 1, 2019, up until May 24, 2022), the evaluation set (from May 25, 2022, to October 31, 2022), and the test set (from November 1, 2022, onward). The reported experimental results are based on the model’s performance on the test set alone.
Appendix B Other Experiment Details
B.1 Length of Retrieval Index Segment of Knowledge Base
Table 5: Length of Retrieval Index Segment of Knowledge Base.
dataset retrieval length metric
mse 4621934.067
4 mae 1303.329
mse 3828479.700
7 mae 1178.574
@@ B.1 Length of Retrieval Index Segment of Knowledge Base
14 mae 1101.172
mse 3747974.421
21 mae 1208.379
mse 3816283.253
28 mae 1243.133
mse 3913793.800
FHSV 35 mae 1332.472
The length of the retrieval segment in a knowledge base significantly affects the accuracy of retrieving the original sequence. Through controlled experiments, we identified the most suitable retrieval segment length for the Fliggy scenario. As shown in Table 5, we incrementally extended the length of the retrieval segment K from 4 units to 35 (corresponding to half to five times the length of the prediction sequence at 7 units), then observed the impact on model performance as measured by MSE and MAE.
The results revealed that during the expansion of the retrieval segment from 4 to 14 units, the predictive performance improved. However, beyond this point, as the retrieval segment continued to lengthen, the predictive performance began to decline. This trend demonstrates that, in practical scenarios, as the retrieval segment grows from short to long, its representation transitions from being information-poor to increasingly mixed and complex.
This situation underscores why, in traditional time series databases, using full-ordered representations for retrieval often yields effective outcomes. By contrast, our K,V design effectively circumvents this issue, providing flexible and accurate retrieval results.
B.2 Effectiveness of Using DTW as Auxiliary Training
In the article, we claimed the effectiveness of using DTW as a feature retrieval technique during the cold start phase to aid in single-epoch training for forecasting models. To validate the necessity of this approach, we designed a set of comparative experiments: a baseline group without DTW-assisted training (named by end to end), and other groups that were each trained with DTW assistance for 1 to 5 epochs. We ensure all experiments had same maximum numbers of training epochs and an identical early stopping strategy. Table 6 shows that the model’s predictive performance is poorest when DTW is not employed; moreover, no better results were obtained when DTW-assisted training exceeded one epoch.
This finding confirms our expectation that, without DTW support during the initial stage, the model struggles to discover high-quality retrieval vectors. Concurrently, our main experiment has previously demonstrated that DTW is not inherently an optimal retrieval scheme geared towards forecasting accuracy. Consequently, prolonging the learning period under DTW guidance can actually restrict the expressive capacity of the forecasting model rather than improving it.
Table 6: Effectiveness of Using DTW as Auxiliary Training.

### v2

@@ 3.4. RACA
3.5. Deployments Procedure
In this section, we briefly introduce the procedure to deploy RATSF in a new field.
3.5.1. TSKB Initialization
Set the initial retrieval length ($L_{r}$), which can be equivalent to the prediction length ($L_{f}$). Set the sequence length for the V sequence ($L_{v}$) to be at least the sum of $L_{r}$ and $L_{f}$. This ensures adequate info for retrieval and forecast. It is suggested to set several $L_{v}$ values, such as $L_{r}$ + $L_{f}$, 2*$L_{f}$ + $L_{r}$, 3*$L_{f}$ + $L_{r}$, etc.,to build several TSKBs for optimal value selection later. Set the rolling window step (stride) to 1 and collect $V$ sequences for each TSKBs. Then, Select the initial $l_{r}$ elements of $V$ sequences to form $K$ sequences.
3.5.2. Identifying Optimal $L_{v}$
Train diverse RATSF models on their respective TSKBs and evaluate each by prediction accuracy.Select the model with the highest prediction accuracy as the optimal RATSF model ($M^{*}$), and identify the corresponding V sequence length ($L_{v}$) as the final chosen value ($L_{v}^{*}$).
3.5.3. Adjusting Retrieval Length $L_{r}$
Adjust the retrieval length ($L_{r}$) to observe the impact on the performance of the optimal model (M*). The adjustment range could be from $0.5L_{f}$ to $L_{v}$, with each adjustment increment being $0.5L_{f}$. Record the $L_{r}$ value that enables M to achieve the highest prediction accuracy, denoted as $L_{r}^{*}$.
3.5.4. TSKB and Model Optimization
With the confirmed optimal retrieval length ($L_{r}*$) and V sequence length ($L_{v}^{*}$), reconfigure the TSKB. Reinitialize and train the RATSF model with the updated TSKB till converge.
4. Experiments
To demonstrate the superiority of our method in service volume forecasting, we first conducted experiments using the Fliggy Hotel Service Volume Dataset (FHSV), showcasing the performance enhancement of typical time-series models when integrated with RATSF. Additionally, to validate the general applicability of our approach, we extended our experiments to three other datasets. Furthermore, through a series of ablation studies, we individually assessed the impact of each key design choice in RATSF to confirm the correctness of our detailed selections.
4.1. Experiment Setup
Datasets. Given that our primary focus lies in the forecasting task within the context of customer service volume management, we initially employed the Fliggy Hotel Service Volume Dataset to substantiate the efficacy of RATSF, the detail description of the dataset is in Appendix A. Additionally, to demonstrate the performance of our approach in other contexts, we employed three classic time-series forecasting datasets: ETT(Zhou et al., 2021), Exchange(Zhao et al., 2019) , Traffic(tra, 2017) .
Models.In Section 3.4, we established that the Retrieval Augmented Cross-Attention (RACA) module is compatible with various Transformer variants. To underscore this versatility, we selected three prominent time-series domain models: Fedformer(Zhou et al., 2022), Autoformer(Wu et al., 2021), and NS-Transformer(Liu et al., 2022), and compared them to the vanilla Transformer. Leveraging the specialized operators within Fedformer and Autoformer that cater to time-series periodicity and trends, such as Autoformer’s seasonal and trend-cyclical initializations, we have seamlessly integrated these models with the RACA module. This integration allows for the effective incorporation of their distinct sequential decomposition capabilities, enhancing the overall processing power. Furthermore, to demonstrate the advantage of our recall mechanism in the RATSF framework, we compared it with ReTime, a RAG-based model that also uses specialized retrieval techniques. Since ReTime has not released their code and their application domain is distinct from our focus on univariate time-series forecasting, we reconstructed the Relational Retrieval method based on their published paper. (It is worth noting that, as of the time of writing this paper, neither MQ-ReTCNN nor ReTime have released their code to the public.)
As shown in Figure 1, the training outcomes of the encoder’s output significantly influence both the retrieval quality and the decoder’s performance. In the main experiments, we will compare the retrieval schemes of RATSF and ReTime. Furthermore, in several subsequent experiments, we will elucidate the differences between RATSF’s retrieval approach and DTW. Consequently, for the majority of the experiments in this paper, to facilitate these comparisons, we have replaced the input to RACA from $\mathbf{X}^{e}_{r}$ to $\mathbf{X}^{b}_{r}$, and we will provide an ablation study to explore the specific effects arising from this substitution.
Forecasting Setting. Like we claim in Section 3.1, we adopt the uni-variant time-series forecasting setting. In the Fliggy’s practical operations, we are required to finalize our forecasting for the upcoming week’s day-by-day staffing needs one week in advance, which is why the actual prediction window period is set at 7 days, represented as 7 tokens. In the travel industry, which Fliggy App’s serving, a quarter generally defines a relatively complete business trend. Therefore, we choose an encoder window length of 98 days—exceeding 90 days and divisible by 14—which consequently results in $\mathbf{X_{o}}$ is 98 units long. And we concate $\left\{x_{t-14},...,x_{t}\right\}$ with a placeholder of length 7 to form $\mathbf{X_{f}}$. To match the length of $\mathbf{X_{f}}$, we set $L_{v}$=21 and $L_{r}$=14. Regarding normalization, we employ unified $\mu$ and $\sigma$ values for whitening operations during the preprocessing stage and inverse normalization is carried out after forecasting. We have applied the same data processing scheme to both the ETT, Exchange and Traffic datasets.
@@ 4.1. Experiment Setup
Table 1. Performance comparison of the Transformer models with their retrieval-based variants on four experimental datasets. ’baseline’ denotes the original model without retrieval, compared to Relational Retrieva from ReTime and our approach RATSF.All results are evaluated using MSE (Mean Squared Error) and MAE (Mean Absolute Error), and the best result is bolded.
@@
Traffic baseline 0.00007 0.00564 0.00006 0.00569 0.00008 0.00643 0.00012 0.00876
Relational Retrieval 0.00007 0.00534 0.00006 0.00512 0.00007 0.00601 0.00009 0.00718
RATSF 0.00006 0.00498 0.00006 0.00472 0.00007 0.00514 0.00007 0.00650
4.2. Main Restult
Table 1 displays the prediction accuracy of the Transformer and its variants on four datasets, comparing performance without RA, with ReTime’s RA, and with our RATSF approach. In the Fliggy Hotel Customer Service Volume Dataset forecasting scenario, the Transformer model’s MAE loss was reduced by 5.34% with ReTime’s Relational Retrieval and by 18% after incorporating RATSF. This improvement could potentially translate to a reduction of approximately 200 in redundant personnel costs in practical management operations. A horizontal analysis of the first row in Table 1 reveals that all compared Transformer-based time-series model variants experienced a significant reduction in MAE after employing Retrieval-Augmented (RA) strategies. Specifically, the adoption of RATSF resulted in respective decreases of 14%, 5%, and 4%. In contrast, the application of ReTime’s retreval method led to improvements of 5.9%, 2.8%, and 1.7% compared to their original designs without RA. These results clearly demonstrate two conclusions: first, effective RA strategies can enhance the performance of time-series Transformer models; second, RATSF exhibits a significant advantage in its retrieval strategy.
@@ 4.2. Main Restult
Additionally, across the ETT, Exchange and the Traffic dataset, the adoption of RATSF design in the Transformer and its time-series optimized variants led to noticeable reductions in MAE metrics. Moreover, among these, the NSTransformer consistently outperforms the other models, likely because its mean-adapted feature proves effective in a broad range of time-series prediction domains.
4.3. Ablation Study
We confirm the effectiveness of several design choices in this section. Section 4.3.1 emphasizes the superiority of RACA’s approach in integrating historical data for improved forecasting. Section 4.3.2 demonstrates the advantages of utilizing model encoder for recalling historical pieces. Section 4.3.3 demonstrates the benefits of incorporating DTW into the training process.
Figure 5. This figure shows RACA’s use of retrieved sequences for a random forecast sequence $\mathbf{X_{f}^{*}}$.The lower half features the top 3 similar sequences of $\mathbf{X_{f}^{*}}$, concatenated along the time axis to form RACA’s input. Attention weights at the first forecast point ($x_{t+1}$) are marked by data bars, with yellow and green indicating RACA’s focus on downward-curving segments. The upper half replicates prediction results (yellow ”pred” lines for $x_{t+1}...x_{t+f}$) and true values (orange ”gt” lines), aligned with the corresponding final $L_{f}$ lengths of the retrieved sequences. This visual comparison highlights the model’s pattern recognition, with RACA’s focus areas correlating to the predictions’ inflection at $x_{t+1}$.For this sample, the model notably focuses on segments indicating ascent and predicts a subsequent increase in value.
Table 2. Ablation of the specifically designed structures within RACA on FHSV scenario.
@@ 4.3. Ablation Study
4.3.1. Integrate Historical Sequences with RACA
@@
A deeper examination disclosed that the Cross Attention mechanism in the RACA design efficiently guides the forecast sequence to concentrate on past trends analogous to the current scenario. In Figure 5, we illustrate how retrieved sequences influence the predicted values for the first forecast point of a specific sequence through RACA’s cross-attention mechanism. The lower half features the top 3 similar sequences, concatenated along the time axis to form RACA’s input. Attention weights at the first forecast point ($x_{t+1}$) are marked by data bars, with yellow and green indicating RACA’s focus on downward-curving segments. The upper half replicates prediction results (yellow ”pred” lines for $x_{t+1}...x_{t+f}$) and true values (orange ”gt” lines), aligned with the corresponding final $L_{f}$ lengths of the retrieved sequences. This visual comparison highlights the model’s pattern recognition, with RACA’s focus areas correlating to the predictions’ inflection at $x_{t+1}$.
Table 3. Comparison of several variants of retrieval representation schemes on the FHSH dataset.
stage methods metric
@@
representation retrieval baseline + encoder mae 1101.172
@@
RATSF with embedding mae 1101.172
mse 3343793.904
decoding RATSF with encoder mae 1004.326
4.3.2. Retrieval Embedding
As previously mentioned, in our main experiments, we aimed to isolate the contribution of the encoder’s output to the retrieval performance. To do this, we utilized the encoder’s output $\mathbf{X}^{e}_{r}$ for the retrieval process and the embedding output $\mathbf{X}^{b}_{r}$ for decoding. In this section, we have completed the discussion of this part of the logic. Here, we will compare different schemes for retrieval and evaluate the effects of various decoding schemes.
We argue that DTW is not the optimal retrieval method and demonstrate this through an experiment on the FHSV dataset, where we compare DTW, a separately trained two-layer MLP, and our current design choice: using the forecasting model’s encoder for retrieval. The upper half of Table 3 presents a comparison of the forecasting precision achieved with these different retrieval embedding strategies. The ’Baseline’ indicates the standard Transformer model without any retrieval mechanism. As the results indicate, the encoder output outperforms the MLP, likely due to the attention mechanism within the Transformer’s encoder, which is adept at identifying and refining the relationships among data points in sequences, thus capturing more accurate patterns.
The lower half of Table 3 compares the decoding performance of RACA when using the Transformer’s embedding output and encoder output as the information source. The experimental results indicate that refining the retrieved sequence information with the Encoder can reduce the MAE by 18.38% to 1004.326, compared to the vanilla Transformer, and by 8.79% compared to using the embedding output. This substantiates the effectiveness of the RATSF architecture demonstrated in Figure 1.
Table 4. Effectiveness of Using DTW as Auxiliary Training.
@@ 4.3.2. Retrieval Embedding
4.3.3. Effectiveness of Using DTW as Auxiliary Training
In Section 3.3.2, we assert the effectiveness of employing DTW as a retrieval method during the cold start phase to facilitate the training of forecasting models for a single epoch. To substantiate the rationale for this strategy, we conducted a comparative experimental study: one baseline group without DTW-assisted training (referred to as ’end-to-end’), and several other groups, each trained with DTW assistance for 1 to 5 epochs. We ensured that all experiments had the same maximum number of training epochs and utilized an identical early stopping strategy. Table 4 demonstrates that the model’s predictive performance is the weakest when DTW is not utilized, as indicated by the highest MSE and MAE values in the last row of Table 4; furthermore, extending DTW-assisted training beyond one epoch did not lead to improved outcomes.
This result aligns with our hypothesis that in the absence of DTW support during the initial training phase, the model has difficulty identifying high-quality retrieval vectors, which are crucial for accurate forecasting. However, as our primary experiments have shown, DTW is not inherently the best retrieval method for enhancing predictive accuracy. Therefore, extending the period of DTW-assisted training may hinder rather than enhance the model’s ability to express and forecast effectively.
5. Conclusion
In conclusion, RATSF stands out in the domain of univariant time-series forecasting by integrating a TSKB that meticulously slices and indexes historical data, leading to superior retrieval precision. This is further enhanced by the Encoder’s representations, which act as a sophisticated retrieval mechanism. Coupled with our innovative RACA mechanism for adeptly merging retrieved segments, we achieve heightened forecast accuracy. The adaptability of our method is highlighted by its compatibility with diverse Transformer variants and its extensive, successful application within Fliggy’s service volume forecasting scenarios.